OpenStudy (anonymous):

for a linear function. f(4)=11 and f(-2)=2. find f(3). i got f(3)=9.5 is that right?

OpenStudy (anonymous):

11-2/4+2=9/6=3/2 this is slope y-2=3/2(x+2) y=3/2x+5 so f(3) = 9/2+5=9/2+10/2=9,5

OpenStudy (anonymous):

thank you